OpenStudy (theraggedydoctor):

If \(x+y=5\) and \(x-y=3\), then \(x^2-y^2=\) 9 15 16 25 34

OpenStudy (turingtest):

Note that \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (theraggedydoctor):

Yes.

OpenStudy (turingtest):

ok, so what's \(x+y\) ?

OpenStudy (theraggedydoctor):

I'm not sure..

OpenStudy (turingtest):

it's given in the problem

OpenStudy (theraggedydoctor):

Oh, 5.

OpenStudy (turingtest):

right what is \(x-y\) ?

OpenStudy (theraggedydoctor):

3

OpenStudy (turingtest):

good now, given that \((x-y)(x+y)=x^2-y^2\), what is \(x^2-y^2\) ?

OpenStudy (theraggedydoctor):

So, \(5 * 3\) = \(5^2-3^2\)=25-9=16

OpenStudy (turingtest):

no, on the left side you substituted correctly, then you decided to make x=5 and y=3 for some reason

OpenStudy (theraggedydoctor):

So just, \(5*3\)

OpenStudy (turingtest):

yep